OpenStudy (anonymous):

Which of the following is the solution to the equation 8^y + 2) = 2^ y?

OpenStudy (anonymous):

a) y = 1 b) y= 3 c) y= -3 d) y= -1

OpenStudy (anonymous):

What you can do is plugin each of the possible answers... But I will show you how to solve.

OpenStudy (anonymous):

First, you need to distribute the 8: 8(y + 2) = 8y + 16 Ok then lets try your answer and see what we get.

OpenStudy (anonymous):

ok then we are stuck with the same problem. Why dont we try inserting the variable and then add before we distribute.

OpenStudy (anonymous):

8(y + 2) What do you want to try first?

OpenStudy (anonymous):

distribute?

OpenStudy (phi):

you start with 8(y + 2) = 2y britanny has the right idea: distribute the 8 8y + 16 = 2y now add -2y to both sides 8y -2y + 16 = 2y + -2y simplify 6y + 16 = 0 add -16 to both sides 6y = 16 divide both sides by 6 y = 16/6 = 8/3 are you sure you typed in the correct question ?

OpenStudy (anonymous):

now look i added ^ those thangs sorry.

OpenStudy (phi):

yes, that makes more sense \[ 8^{y+2} = 2^y \] the trick is to know that 8= 2*2*2 , or 2^3 \[ \left(2^3\right)^{y+2} = 2^y \] use the rule that \[ \left(x^a\right)^b= x^{ab} \] to rewrite the problem as \[ 2^{3(y+2)} = 2^y \] now set the exponents equal to each other 3(y+2) = y can you solve this ?

OpenStudy (anonymous):

no

OpenStudy (phi):

3(y+2) = y distribute the 3 on the left side.

OpenStudy (anonymous):

3y+6=y ?

OpenStudy (phi):

yes, now add -y to both sides: 3y+ -y + 6=y + -y can you simplify that ?

OpenStudy (anonymous):

3+6=y ? idk that one confused me

OpenStudy (phi):

3y+ -y + 6=y + -y on the right side you have y + -y which says " 1y take away 1 y" (you get zero) on the left side you have 3 y and take away 1 y can you do it now ?

OpenStudy (anonymous):

3y+6=y ?

OpenStudy (phi):

that is what you start with. now add -y to both sides 3y -y + 6 = y -y what is y-y ?

OpenStudy (anonymous):

y-y is nothing? just 0

OpenStudy (phi):

yes, so 3y -y + 6 = y -y simplifies (on the right side) to 3y - y + 6 = 0 now what is 3 y's take away 1 y ?

OpenStudy (anonymous):

2y's

OpenStudy (phi):

so now you have 2y + 6 =0 next, add -6 to both sides. can you do that ?

OpenStudy (phi):

2y + 6 -6 = 0-6 or 2y = -6 now divide both sides by 2

OpenStudy (anonymous):

-3 is the answer??

OpenStudy (phi):

yes

OpenStudy (anonymous):

thanks